# Critical Phase Transition in a Large Language Model

### v1

@@ Abstract
The performance of large language models (LLMs) strongly depends on the temperature parameter. Empirically, at very low temperatures, LLMs generate sentences with clear repetitive structures, while at very high temperatures, generated sentences are often incomprehensible. In this study, using GPT-2, we numerically demonstrate that the difference between the two regimes is not just a smooth change but a phase transition with singular, divergent statistical quantities. Our extensive analysis shows that critical behaviors, such as a power-law decay of correlation in a text, emerge in the LLM at the transition temperature as well as in a natural language dataset. We also discuss that several statistical quantities characterizing the criticality should be useful to evaluate the performance of LLMs.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive performance in various tasks, including machine translation and code generation. To optimize their performance, tuning the hyperparameters is crucial. An example of such hyperparameters is the temperature $T$. This parameter scales the logits of tokens to control the sharpness of distribution of the next tokens. Decreasing temperature makes the probability measures concentrate on the most likely token, while increasing temperature flattens the distribution. Empirically, at very low temperatures, generated sentences are less diverse with clear repetitive structures. Conversely, generated sentences are often disordered and incomprehensible at very high temperatures. These observations raise the question of whether this change induced by the temperature is quantitative or qualitative.
Figure 1: Schematic pictures of critical phenomena in physics and their analogies: (A) Phase transition in a ferromagnetic Ising model, showing how magnetization exhibits a singularity as a function of temperature in the infinite system size limit. This singular point, which is called the phase transition point, separates the parameter space into ordered and disordered phases. (B) An analogous phase transition in GPT-2. The integrated correlation of POS sequences generated by GPT-2 diverges below the phase transition point $T_{c}$ and converges to a finite value above the point. At $T_{c}$, critical phenomena are observed. (C) A conjectured relation between natural languages, GPT-2, and their criticality within a parameter space where each element represents a distribution of sequences.
The problem of defining qualitative changes without subjectivity, such as arbitrary thresholds, is itself a major challenge. A well-defined strategy to address this question is found in the literature of phase transitions studied in statistical physics. Phase transitions are characterized by singularities in physical quantities. For example, in a ferromagnetic Ising model with a finite number of spins, the degree to which the spins are aligned, known as magnetization, decreases smoothly with increasing temperature. As the number of spins increases, this change becomes sharper and, in the infinite system size limit, it eventually becomes a strictly singular change at the phase transition temperature (Domb, 2000). This transition temperature separates the parameter space unambiguously into two phases, that is, the ordered and the disordered phases (see Fig. 1 (A)).
Phase transitions accompanying critical phenomena, e.g., power-law decays and diverging scales in the correlation functions, are said to be critical. This is in contrast with long- and short-range correlations, both of which rapidly decay to the limiting value with finite scales. At the critical point, the divergent scales of correlations allow for complex behaviors that are neither ordered nor disordered. Such behaviors have been discussed in various scientific disciplines, including computer science, biology, finance, and others (Bak et al., 1988; Langton, 1990; Munoz, 2018). Natural language is also a complex object in the sense that it is constrained by rules such as syntax and yet has infinite variety, colloquially referred to as a critical phenomenon due to a power-law correlation decay (Li, 1989; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023). However, this remains speculative and it is even unclear what kind of phase transition should be behind the putative criticality, as natural language does not have a controllable parameter such as temperature.
@@
In this work, we use GPT-2 (Radford et al., 2019) as an effective model of natural language to systematically study the effects of the temperature parameter. In our analysis, we numerically calculate statistical properties of part-of-speech (POS) sequences $(y_{0},\cdots,y_{N-1})$ generated by GPT-2. Throughout this paper, we interpret index $t$ as time and a POS sequence as a time series 111 We might also consider $t$ as a spatial position. Under this interpretation, the sequence can be regarded as a one-dimensional configuration of a Potts spin system with $N$ interacting spins (Wu, 1982). The integrated correlation $\tau$ in Sec. 3.1 corresponds to a static susceptibility, the power spectrum $S$ in Sec. 3.2 to the structure factor, and the vector $\{v_{a}(t)\}$ in Sec. 3.3 to the magnetization of each spin. . Note that causality is broken in the step of POS tagging, since the POS tag assigned to a word can be affected by the subsequent context. Our extensive numerical analyses on the statistical quantities strongly suggest that a phase transition occurs at temperature $T_{c}\approx 1$ as illustrated in Fig. 1 (B). Below $T_{c}$, sequences have long-range correlation and repetitive structures. On the other hand, above $T_{c}$, the correlation converges to zero, and repetitive structures disappear. We further demonstrate that sequences at the transition point are critical, with a critical decay in a correlation and dynamical slowing down. The observed behaviors exhibit several intriguing features that are not known in equilibrium statistical-mechanical systems. At high temperatures, we find that the system is not simply disordered even though the correlation decays faster than the one at the critical point. The low-temperature phase is ordered, with characteristic power spectra implying its chaotic nature. Finally, we measure the same statistical quantities for a natural language dataset, providing the interpretation that natural languages are associated with the critical point of GPT-2 (see Fig. 1 (C)).
@@
  • We systematically investigate statistical quantities of sequences generated by GPT-2 at various sampling temperatures to clearly characterize the effect of temperature and suggest the possibility of the use of statistical measures for evaluating LLMs.
  • Our work provides the first convincing numerical evidence that a practical LLM exhibits a phase transition. This finding introduces novel approaches to understanding the mechanism of LLMs from a perspective of critical phenomena studied in statistical physics.
  • We show that the criticality of GPT-2 is intriguing also from a statistical-mechanics perspective, which may be caused by the complex architecture of LLMs with vast numbers of parameters trained on large-scale corpora. We also discuss the peculiar temporal structures at high and low temperatures.
  • We conduct a numerical analysis of the statistical properties of a natural language dataset and discuss connections with the criticality of GPT-2.
@@ 1.1 Related Work
Phase transitions in practical language models: The performance of LLMs sometimes improves in unexpected manners with increasing model scales (Brown et al., 2020; Srivastava et al., 2022; Ganguli et al., 2022). Wei et al. (2022) named “abilities that are not present in smaller-scale models but are present in large-scale models” emergent abilities, mentioning phase transitions, although Schaeffer et al. (2024) has argued that such emergent abilities may not be essentially qualitative changes. Grokking (Power et al., 2022) in neural networks, the emergence of generalization after overfitting with increasing learning steps, is also sometimes discussed in connection with phase transitions (Liu et al., 2022a; Thilak et al., 2022; Nanda et al., 2023; Varma et al., 2023). Liu et al. (2022b) and Zhu et al. (2024a) reproduced grokking with simple language models and tasks. However, emergent abilities and grokking in practical language models have not yet been formally characterized as phase transitions, as discussed in our study. It is also important to note that these phenomena are changes in performance induced by model scale or learning steps, contrary to the changes in generated sequences induced by sampling temperature on which we focus.
@@
During the preparation of our manuscript, we noticed an independent work by Bahamondes (2023), which measured statistical quantities of sequences of tokens or embedding vectors generated by GPT-2 trained on OpenWebText dataset (Wolfram Research, 2019; Radford et al., ). This study claimed that a phase transition occurs at $T\approx 0.1$. However, they did not mention the dependence of correlation on distance or the asymptotic behavior in the limit of infinite sequence length. We should also note that the phase transition discussed in Bahamondes (2023) would be distinct from the one we study in this work, as the temperature scale is very different.
Theoretical analyses of phase transitions in language models: Some theoretical studies have examined phase transitions in mathematical models to explain phenomena such as emergent abilities (Chang, 2023) and grokking (Žunkovič and Ilievski, 2022; Rubin et al., 2024), although some of them are not necessarily limited to LLMs. Cui et al. (2024) demonstrated that a simple model with a self-attention mechanism exhibits a phase transition between phases corresponding to positional and semantic mechanisms. However, all of them have discussed different phenomena from a sampling-temperature-induced phase transition in generated sequences, which we discuss in this study.
@@
Correlation in natural languages: Power-law decays in correlations and mutual information have been observed across different natural language datasets222 Note that the earlier studies often referred to correlations decaying in a power-law function as long-range correlations. In this study, by long-range correlations we mean correlations that converge to a non-zero value. (Li, 1989; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023). These studies, however, did not measure the correlation or mutual information between POS tags, which is our focus. Futrell and Levy (2017) and Futrell et al. (2019) have measured mutual information between POS tags, only to show that the mutual information decreases with distance.
Correlation in texts generated by language models: The analysis of correlation and other statistical quantities in texts generated by language models have been conducted numerically and theoretically by Takahashi and Tanaka-Ishii (2017); Shen (2019); Takahashi and Tanaka-Ishii (2019), and Mikhaylovskiy and Churilov (2023), which concluded that the presence of power-law decays in correlations can depend on models, methods, and setups. These studies have not examined how the correlations depend on parameters such as temperatures. Lippi et al. (2019) has found a power-law decay in a correlation in texts generated by language models with varying temperatures in comparison with natural language datasets. However, the existence of phase transitions was not discussed in their works.
@@ 2 Setup
We analyze sequences generated by GPT-2 (Radford et al., 2019) using Hugging Face transformers library (Wolf et al., 2020), with temperature sampling (Ackley et al., 1985). Each sequence starts with the single beginning-of-sentence token and ends once the end-of-sentence token appears or when its length reaches the default length limit of $1024$. The $t$-th token $x_{t}$ is sampled according to the softmax distribution $P(x_{t}|x_{0},\cdots,x_{t-1})\propto\exp({-H(x_{t}|x_{0},\cdots,x_{t-1})/T})$ without top-$k$ (Fan et al., 2018) or top-$p$ (Holtzman et al., 2020) sampling strategies. $H(x_{t}|x_{0},\cdots,x_{t-1})$ is the logit of $x_{t}$, while $T$ is the temperature parameter, although it does not necessarily coincide with physical temperature. We then tokenize each of the generated sequences using the GPT-2 tokenizer. To estimate statistical quantities precisely, it is useful to map each generated text into a sequence of variables that take a small number of states while preserving linguistic information. For this purpose, we employ POS tagging using the spaCy library (Montani et al., 2023) with en_core_web_sm pipeline. This process transforms the generated texts into POS sequences $(y_{0},\cdots,y_{N_{\text{POS}}-1})$, where $y_{t}$ takes one of 18 different POS tags, that is, 17 Universal POS tags plus SPACE.
We sampled $3.2\times 10^{5}$ POS sequences at each temperature. Sampling at any given temperature took less than a day using a single NVIDIA A100. To study the POS sequences of length $N$, we have used only samples that are longer than $N$ and extracted the first $N$ tags $(y_{0},\cdots,y_{N-1})$ from the beginning. For example, a POS sequence of $N_{\text{POS}}=300$ is used for $N=256$ but not for $N=512$. The number of samples to compute statistical quantities is $1.6\times 10^{5}$. We have confirmed that this sampling method does not qualitatively affect the results presented below (see App. A). Error bars in the figures represent an $80\%$ confidence interval estimated using the symmetric bootstrap-$t$ method (Hall, 1988).
@@ 3 Numerical results for GPT-2 sequences
In this section, we conduct a detailed statistical analysis of sequences generated by GPT-2 to understand the effect of temperature on their structure and dynamics. More specifically, we study the correlation between POS tags, the power spectra of these sequences, and the time evolution of POS tag distributions.

### v2

# Critical Phase Transition in Large Language Models

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive performance. To understand their behaviors, we need to consider the fact that LLMs sometimes show qualitative changes. The natural world also presents such changes called phase transitions, which are defined by singular, divergent statistical quantities. Therefore, an intriguing question is whether qualitative changes in LLMs are phase transitions. In this work, we have conducted extensive analysis on texts generated by LLMs and suggested that a phase transition occurs in LLMs when varying the temperature parameter. Specifically, statistical quantities have divergent properties just at the point between the low-temperature regime, where LLMs generate sentences with clear repetitive structures, and the high-temperature regime, where generated sentences are often incomprehensible. In addition, critical behaviors near the phase transition point, such as a power-law decay of correlation and slow convergence toward the stationary state, are similar to those in natural languages. Our results suggest a meaningful analogy between LLMs and natural phenomena.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive performance in various tasks, including machine translation and code generation. Uncovering the origin of the unexpectedly good performance of LLMs is essential both from theoretical and practical perspectives, as it would allow us to devise even better LLMs. A complete theory for LLMs is yet to be developed, but it definitely needs to take into account the fact that LLMs sometimes show qualitative changes when varying parameters. Indeed, various qualitative changes play a crucial role in LLMs. For example, it is empirically known that LLMs can exhibit emergent abilities: The performance of LLMs sometimes improves unexpectedly when increasing model scales (Brown et al., 2020; Srivastava et al., 2022; Ganguli et al., 2022; Wei et al., 2022), although some argues that the emergent ability may be an artifact due to the arbitrary choice of metrics (Schaeffer et al., 2024). The grokking phenomenon (Power et al., 2022), where, during the learning process, the generalization loss drastically decreases long after the convergence of the training loss, is also an example of such qualitative changes that are observed in language models (Liu et al., 2022b; Zhu et al., 2024a).
Qualitative changes are ubiquitous in nature, often expressing themselves as phase transitions. For instance, as temperature increases, ferromagnets undergo a phase transition from the ordered phase, in which the magnetic spins are aligned and the materials are magnetized, to the disordered phase with zero net magnetization. Interestingly, the phase transition point that separates these two phases is rigorously defined by singular, divergent statistical quantities in the limit where the system size is infinitely large, see Fig. 1 (A). This point does not depend on any subjective factors, such as metrics and thresholds.
These raise an intriguing question: Are qualitative changes observed in LLMs phase transitions? If so, it suggests a meaningful analogy between LLMs and natural phenomena and opens the possibility of understanding LLMs through theories and methods for studying phase transitions in the natural world. In this study, we numerically investigate the possibility of phase transitions in LLMs. We shall first focus on the change in the statistics of generated texts with the temperature parameter. This parameter is easy to control and known to induce an evident change to generated texts: At low temperatures, clear repetitive structures appear in the texts, whereas, at very high temperatures, they are often incomprehensible. This behavior reminds us of phase transitions between a periodic, ordered phase and a disordered phase when changing physical temperature. Another motivation is the fact that the decay of correlation in natural languages follows a power law (Li, 1989; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023). The power-law decay of correlation is one of the important features of critical phenomena, which occur near the phase transition point between the ordered and disordered phases. These suggest the existence of a phase transition in LLMs, as illustrated in Fig. 1 (B): The phase transition occurs at a certain temperature $T_{c}$, at which the LLM exhibits critical behaviors similar to those in natural languages.
@@
To examine the above speculation, we have conducted extensive statistical analysis of sequences generated by LLMs, to detect phase transitions that would exist in the limit of infinitely long sequences. Our numerical results strongly suggest that a phase transition occurs at temperature $T_{c}\approx 1$. Below $T_{c}$, sequences have long-range correlation and repetitive structures. On the other hand, above $T_{c}$, the correlation converges to zero, and repetitive structures disappear. We further demonstrate that sequences at the transition point have critical properties, such as a power-law decay in a correlation and slow convergence toward the stationary state. The observed behaviors exhibit several intriguing features that are not known in equilibrium statistical-mechanical systems; at high temperatures, we find that the system is not simply disordered even though the correlation decays faster than the one at the critical point. The low-temperature phase is ordered, with characteristic power spectra implying its chaotic nature. Finally, we show natural language datasets have critical properties as well as the LLMs at the transition temperature, implying that the scenario presented in Fig. 1 (B) indeed holds.
Figure 1: Schematic pictures of phase transitions and critical phenomena in physics and LLMs: (A) Phase transition in a ferromagnetic Ising model, showing how susceptibility exhibits a singularity as a function of temperature in the infinite system size limit. This singular point, which is called the phase transition point, separates the parameter space into ordered and disordered phases. (B) A conjectured relation between LLMs and natural languages within a parameter space where each element represents a distribution of sequences.
@@
  • We propose to formulate qualitative changes in LLMs as phase transitions studied in statistical physics, providing the first strong evidence that practical LLMs exhibit a phase transition.
  • We conduct a numerical analysis of the statistical properties of natural language datasets and discuss connections with the criticality of LLMs.
  • We show that temporal structures of texts generated at high and low temperatures are peculiar from a statistical-physics perspective, which may be caused by the complex architecture of LLMs with vast numbers of parameters trained on large-scale corpora.
@@ 1.1 Related Work
Phase transitions in language models: Emergent abilities and grokking are discussed in connection with phase transitions (Wei et al., 2022; Liu et al., 2022a; Thilak et al., 2022; Nanda et al., 2023; Varma et al., 2023). In these studies, their formulation of phase transitions is not rigorous. Indeed, Schaeffer et al. (2024) has argued that emergent abilities may not be actual qualitative changes. Some theoretical studies formulate emergent abilities, grokking, and other phenomena in language models as phase transitions, yet they consider only mathematical models, not practical LLMs (Žunkovič and Ilievski, 2022; Chang, 2023; Rubin et al., 2024; Cui et al., 2024). Contrary to these earlier studies, we show that a phase transition occurs in practical LLMs. We also note that the earlier studies do not focus on the temperature-induced change in generated texts.
@@
During the preparation of our manuscript, we noticed an independent work by Bahamondes (2023), which measured statistical quantities of sequences of tokens or embedding vectors generated by GPT-2 trained on OpenWebText dataset (Wolfram Research, 2019; Radford et al., ). This study claimed that a phase transition occurs at $T\approx 0.1$. However, they did not mention the dependence of correlation on time interval or the asymptotic behavior in the limit of infinite sequence length. We should also note that the phase transition discussed in Bahamondes (2023) would be distinct from the one we study in this work, as the temperature scale is very different.
@@
Correlation in natural languages and generated texts: Power-law decays in correlations and mutual information have been observed across different natural language datasets111 Note that the earlier studies often referred to correlations decaying in a power-law function as long-range correlations. In this study, by long-range correlations we mean correlations that converge to a non-zero value. (Li, 1989; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023). These studies, however, did not measure the correlation or mutual information between part-of-speech (POS) tags, which is our focus. Futrell and Levy (2017) and Futrell et al. (2019) measured mutual information between POS tags, yet they did not show the power-law decay. The analysis of correlation and other statistical quantities in texts generated by language models have also been conducted numerically and theoretically in earlier studies (Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Mikhaylovskiy and Churilov, 2023), which concluded that the presence of power-law decays in correlations can depend on models, methods, and setups. These studies have not examined how the correlations depend on parameters such as temperature. Lippi et al. (2019) has studied a power-law decay in a correlation in generated texts with varying temperatures in comparison with natural language datasets. However, the existence of phase transitions was not discussed in their works.
Criticality in biological systems: Critical behaviors have also been observed in various biological systems, including neural activity (Munoz, 2018). Previous research has shown that Ising models fitted to neural activity exhibit a phase transition near the temperature where the models are close to the data (Tkacik et al., 2006; Mora and Bialek, 2011). It is noteworthy that their models and data are quite small and simple compared to LLMs.
2 Phase transitions and critical phenomena
In statistical physics, phase transitions are defined by the singularity of a statistical quantity. To be more precise, when any quantity depending on the system changes with a singularity at a parameter value, there exists a phase transition at the point. We stress that true singularity can exist only in the large size limit, where the degrees of freedom of the system is infinitely large.
Near the phase transition point, the system can exhibit critical phenomena. In this case, the phase transition point is also called the critical point. The essential characteristic of critical phenomena is the divergence of the length and time scales of the system, i.e., the scale-free correlation that follows a power law function. This divergence of the scale leads to the divergent integrated correlation, the slowing down of dynamics, and other intriguing phenomena. These phenomena are not observed in ordered or disordered phases, where the correlation is long- or short-range, decaying rapidly to the limiting value with finite scales. Critical behaviors have been discussed in various scientific disciplines, including computer science, biology, finance, and others (Bak et al., 1988; Langton, 1990; Munoz, 2018; Bouchaud, 2024).
In general, theoretically proving the existence of a phase transition is so challenging that it has only been achieved in a few simple theoretical models because we need to consider the large system size limit. However, by estimating statistical quantities with varying system sizes, we can follow the dependence of the quantities on the size to discuss how the quantities diverge. For example, in a ferromagnetic Ising model with a finite number of spins, the degree of response of spins to the magnetic field, known as susceptibility, increases smoothly near the critical temperature. As the system size, i.e., the number of spins, increases, this change becomes sharper and, in the large size limit, it eventually becomes a strictly singular change at the critical point (Domb, 2000), as illustrated in Fig. 1 (A). This transition temperature unambiguously separates the parameter space into two phases, that is, the ordered and the disordered phases. Our approach in this study also follows this method.
@@ 3 Setup
We generate texts using pretrained LLMs with temperature sampling (Ackley et al., 1985). Each sequence starts with the single beginning-of-sentence token and ends once the end-of-sentence token appears or when its length reaches the default length limit. The $t$-th token $x_{t}$ is sampled according to the softmax distribution $P(x_{t}|x_{0},\cdots,x_{t-1})\propto\exp({-H(x_{t}|x_{0},\cdots,x_{t-1})/T})$ without top-$k$ (Fan et al., 2018) or top-$p$ (Holtzman et al., 2020) sampling strategies. $H(x_{t}|x_{0},\cdots,x_{t-1})$ is the logit of $x_{t}$, while $T$ is the temperature parameter, although it does not necessarily coincide with physical temperature. We then tokenize each of the generated sequences.
To estimate statistical quantities precisely, it is necessary to map each generated text into a sequence of variables that take a small number of states while preserving linguistic information. For this purpose, we map the text to the sequence of universal POS tags. This mapping is useful as it can be easily applied to different languages. Note that our analyses in the following part of this work are applicable to another mapping, such as one-to-semantic tags. In addition, the phase transition is observed with the mapping based on characters, as we will show in App. I. This mapping transforms the generated texts into POS sequences $(y_{0},\cdots,y_{N_{\text{POS}}-1})$, where $y_{t}$ takes one of 18 different POS tags, that is, 17 Universal POS tags plus SPACE. To study the POS sequences of length $N$, we choose POS sequences that are longer than $N$ from all the samples, and analyze the first $N$ tags $(y_{0},\cdots,y_{N-1})$ from the beginning for each sequence. For example, a POS sequence of $N_{\text{POS}}=300$ is used for $N=256$ but not for $N=512$. We have confirmed that this sampling method does not qualitatively affect the results presented below, as discussed in App. A. Throughout this paper, we interpret index $t$ as time and a POS sequence $(y_{0},\cdots,y_{N-1})$ as a time series. In this case, the length $N$ of the time series corresponds to the system size 222 Strictly speaking, causality is broken in the step of POS tagging, i.e., the POS tag assigned to a word can be affected by the subsequent context, yet this effect should not be essential. Indeed, we can observe the phase transition with mapping based on characters in App. I, where causality is not broken. .
Since extensive sampling is needed for precise statistical analysis, we used GPT-2 small with 124M parameters (Radford et al., 2019), using Hugging Face transformers library (Wolf et al., 2020) (We also show the results for Japanese GPT-2 medium with 361M parameters (Zhao and Sawada, ; Sawada et al., 2024) in App. H and for Pythia with 70M parameters (Biderman et al., 2023) in App. I). For POS tagging, we use the spaCy library (Montani et al., 2023) with en_core_web_sm pipeline. We sampled $3.2\times 10^{5}$ POS sequences at each temperature. Sampling at any given temperature took about 17 hours using a single NVIDIA A100. The number of samples to compute statistical quantities is $1.6\times 10^{5}$, since not all sequences are longer than $N$. Error bars in the figures represent an $80\%$ confidence interval estimated using the symmetric bootstrap-$t$ method (Hall, 1988) [^3] We attach the codes as supplemental material..
@@ 4 Numerical results for GPT-2 sequences
In this section, we present a detailed statistical analysis of sequences generated by GPT-2 to understand the effect of temperature on their structure and dynamics. More specifically, we study the correlation between POS tags, the power spectra of sequences, and the time evolution of POS tag distributions.